OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

x=-1. You can get this if you use the elimination method, or substitution method.

OpenStudy (anonymous):

solve for y in the first equation and get y=-3+x now plug into the 2nd equation x+3(-3+x) = 5 x-9+3x=5 4x-9 = 5 4x=14 x= 14/4

OpenStudy (anonymous):

and y=2, again, this can be achieved by elimination or substitution method.

OpenStudy (anonymous):

oops made a mistake solving for y in the first equation gives you y = x+3 and now plug that into 2nd equation x+3(x+3) = 5 x+3x+9 = 5 4x +9 = 5 4x=-4 x= -1

OpenStudy (anonymous):

x= 2x^2 + 4 / 5^2 - 264 + 3y / 7